If tan A * tan B = ½, then (5 – 3 cos 2A)(5 – 3 cos 2B) = 

Grade:11

1 Answers

ANSHUMAN SINGH
5 years ago
(5-3cos2a)(5-3cos2b)
=(5-3(1-tan2a/1+tan2a) ) ×(5-3(1-tan2b/1+tan2b))
=(2+8tan2a/1+tan2a) × (2+8tan2b/1+tan2b)
=4+16(1+tan2a +tan2b) / (1+tan2a +tan2b) + 0.25
Now take 16 common in numerator , then terms will cancel out and answer is 16.